OpenStudy (anonymous):

Paul is a salesperson who earns both a salary and a commission on units sold. In one week, he sold 25 units and earned $1,200. In another week, he sold 20 units and earned $1,075. Which of the following slope-intercept form equations models Paul’s weekly pay in terms of units sold? A. y=25x+575 B. y = 25x + 1,200 C. y=-1/25x-68 D. y=1/25x-23

OpenStudy (anonymous):

@ashley_f97 you think you know it?

OpenStudy (ashley_f97):

I'm trying to figure it out

OpenStudy (anonymous):

okie dokie

OpenStudy (mathstudent55):

The salary is a one-time fixed amount. In this case it's unknown; let's call it s. The commission is paid out per unit sold. Let's call it c. If he sells 25 units, he makes 25c dollars in comminssion plus the fixed s dollars. In one week he earned $1200, so 25c + s = 1200 The other week, he sold 20 units earning 20c commission and earning a total salary of $1075, so 20c + s = 1075. We have: 25c + s = 1200 20c + s = 1075 If you solve this system of equations, you will know how much his fixed salary is and how much his commission is. Then you can write a function to find his weekly salary and comapre with the choices.

OpenStudy (anonymous):

@ashley_f97 you get it friend?

OpenStudy (ashley_f97):

Yea I get it

OpenStudy (anonymous):

I got it its a

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

welcome to 50

OpenStudy (mathstudent55):

Correct. A is it.

OpenStudy (anonymous):

thanks @mathstudent55

OpenStudy (ashley_f97):

lol thanks

OpenStudy (mathstudent55):

yw